# Learning In-Hand Translation Using Tactile Skin With Shear and Normal Force Sensing

### v1

@@ 6.2 Three-Axis Tactile Sensing Enables Superior Out-of-Domain Adaptation
6.3 Experimental Analysis
Latent Space Analysis. We analyze the extrinsics vector from our experiments in Figure 7A with t-SNE to produce 3D plots in Figure 7B. The plots visualize from the whole duration of 15 rollouts per policy (5 per object). We find a correlation between high dispersion of clusters and complete policy failure. U3-Axis completely failed with the hammer at 20 degrees and Proprio Only completely failed with the water bottle at 5 degrees. The clusters for these failed rollouts are more dispersed compared to more successful rollouts. In contrast, the S3-Axis policy was more successful overall and exhibited a dense cluster of from all rollouts.
Tactile Policies Explore More Finger Gaits. We analyze the joint states of all policies from our experiments in Figure 7A. We compare the standard deviation of intersections through Poincaré sections of the phase portraits for each finger motor and find that on average, all tactile policies explore more joint states relative to Proprio Only, particularly for the hammer and water bottle experiments. Figure 7C shows the contrast in joint state exploration in phase portraits of S3-Axis and Proprio-Only, where S3-Axis has an increase of 35% standard deviation of intersections through the Poincaŕe section. Greater joint state exploration is potentially a factor in task success and an indicator of gait adaptation; however, other variables such as gait cycle timing and finger coordination are also important to consider. More analysis is in the Appendix.
Real-World Videos. We show videos of real-world experiments with our methods and ablations on our Project Website. These videos can be used to observe task metrics and policy gait adaptation.
7 Conclusion
@@
Limitations and Future Work. We mainly study the new sensor model for touch simulation and learning, and do not consider vision as an input. We will need vision for precise goal-specification. Our control policy is frozen during deployment, but it would be interesting to use real world tactile feedback to fine tune the policy. We can also investigate domain adaptation methods and sensor simulation improvements for continuous tactile signals, which could enable better performance.
Acknowledgments
@@
This work was conducted while Jessica Yin was an intern and Haozhi Qi was a research fellow with Meta FAIR. In their academic roles at UC Berkeley, Haozhi Qi and Jitendra Malik are supported in part by ONR MURI N0001421-1-2801. In her academic role at the University of Pennsylvania, Jessica Yin was supported by the NSF Graduate Research Fellowship under Grant No. 202095381. We thank Mustafa Mukadam, Mike Lambeta, Tingfan Wu, Luis Pineda, Taosha Fan, Patrick Lancaster, Mrinal Kalakrishnan, Raunaq Bhirangi, Carolina Higuera, Akash Sharma, Suddhu Suresh, and William Yang for helpful discussions and feedback throughout this work. We thank Dr. Nadia Figueroa, Ho Jin Choi, Tianyu Li, and Harshil Parekh for access and assistance with the Franka robot arm and Optitrack system for hardware experiments.
Appendix A Demonstrations
A.1 Generalization to Diverse Objects
We demonstrate the generalization of our 3-axis tactile policies to diverse objects, outside of the 4 real world objects in our test set. Please see the supplemental video for footage.
A.2 Real ReSkin and U3-Axis Model Comparison
Here, we compare our U3-Axis tactile skin model to a representative example of real ReSkin. This is the same taxel and same rollout as the examples shown in Figure 3.
Appendix B Experiments
B.1 Object Dataset
1.
Cylinder (ID). This 3D printed uniform cylinder evaluates the policy on an ID object. It has a cm diameter and cm length, weighing g.
2.
Hammer (OOD). The hammer tests adaptation to a skewed COM and rectangular handle geometry. It is 3D printed and inspired by a real hammer from McMaster-Carr111https://www.mcmaster.com/5914A15/. All dimensions are scaled by 200% to match the scale of an Allegro Hand compared to a human hand. It has a total length of cm, and the handle is cm by cm. The hammer head is cm by cm by cm. The total weight is g.
3.
Screwdriver (OOD). The screwdriver evaluates adaptation to complex geometry and multiple contacts. It is 3D printed and inspired by a real screwdriver from McMaster-Carr222https://www.mcmaster.com/7127A34/, also scaled by 200%. The handle diameters range from cm to cm, and the handle is cm. The total length is cm, and it weighs g.
4.
Water Bottle (OOD). The water bottle tests adaptation to a variable COM and non-uniform cylindrical geometry. The water bottle diameter ranges from to cm, and the total weight is g. It contains g of water.
B.2 Real World Evaluation Results
Summary tables for all real world experiments are reported below.
| Water Bottle - Variable COM | ||||
| Hand Tilt | Policy | Success (of 5) | Avg. Dist. (cm) | Avg. Vel. (cm/s) |
| 0 deg. | Signed 3-Axis | |||
| Unsigned 3-Axis | ||||
| Signed Shear Only | ||||
| Normal Only | ||||
| Proprioception Only | ||||
| 5 deg. | Signed 3-Axis | |||
| Unsigned 3-Axis | ||||
| Signed Shear Only | ||||
| Normal Only | ||||
| Proprioception Only | ||||
| 10 deg. | Signed 3-Axis | |||
| Unsigned 3-Axis | 0 | 0 | ||
| Signed Shear Only | 0 | 0 | ||
| Normal Only | 0 | 0 | ||
| Proprioception Only | 0 | 0 | ||
| Cylinder - ID Object | ||||
| Hand Tilt | Policy | Success (of 5) | Avg. Dist. (cm) | Avg. Vel. (cm/s) |
| 0 deg. | Signed 3-Axis | |||
| Proprioception Only | ||||
| 5 deg. | Signed 3-Axis | |||
| Proprioception Only | ||||
| 10 deg. | Signed 3-Axis | |||
| Proprioception Only | ||||
| 15 deg. | Signed 3-Axis | |||
| Proprioception Only | ||||
| Hammer - Skewed COM | ||||
| Hand Tilt | Policy | Success (of 5) | Avg. Dist. (cm) | Avg. Vel. (cm/s) |
| 15 deg. | Signed 3-Axis | |||
| Unsigned 3-Axis | ||||
| Signed Shear Only | ||||
| Normal Only | ||||
| Proprioception Only | ||||
| 20 deg. | Signed 3-Axis | |||
| Unsigned 3-Axis | ||||
| Signed Shear Only | ||||
| Normal Only | ||||
| Proprioception | ||||
| Screwdriver - Challenging Geometry | ||||
| Hand Tilt | Policy | Success (of 5) | Avg. Dist. (cm) | Avg. Vel. (cm/s) |
| 0 deg. | Signed 3-Axis | |||
| Unsigned 3-Axis | ||||
| Signed Shear Only | ||||
| Normal Only | ||||
| Proprioception Only | ||||
B.3 Gait Analysis
We show the phase portraits of the middle finger motor for all policies evaluated in the Hammer, 20 degree hand tilt experiment. For this specific example, we analyze joint state exploration by measuring the standard deviation of intersections through the Poincarè section, which is a plane through and . Here, S3-Axis has +35% greater deviation, U3-Axis has +18% deviation, Signed Shear Only has +26% deviation, and Normal Only has -40% deviation, relative to Proprio Only. Although Normal Only is the exception here, we find that on average with the highlighted experiments in Figure 7, tactile policies generally explore more joint states.
Appendix C Implementation Details
C.1 Reward Function
This is the reward function we use to train the oracle policy:
The in-hand translation task is defined as , where is the object’s current position and is the object’s goal position. In this work, randomized goal positions are a curriculum to produce a versatile, robust finger gait. However, using only this reward term results in policies that exploit inaccurate sliding contact simulation that do not transfer from sim-to-real. Thus, we add additional penalties on object state. penalizes object rotation, as another reward when the object reaches the goal position, and to penalize dropping the object.
C.2 Hyperparameters
| Reward Function | |
| Reward Term | Scale, |
| Simulation Physics | |
| Parameter | Randomization Range or Value |
| Object Scales | |
| Object Mass | kg |
| Object Center of Mass (COM) | m |
| Coefficient of Friction | |
| PD Controller Stiffness | |
| PD Controller Damping | |
| Contact Offset | |
| Bounce Threshold Velocity | m/s |
| Max Depenetration Velocity | m/s |
| Tactile Skin Simulation | |
| Term | Value |
| Sensing Range Threshold (Beyond Taxel) | cm |
| Sensing Range Randomization | |
| Sensor Noise | |
| Threshold for Binary Values | N |
| Binary Filter for Real ReSkin | |||||
| Axis | Threshold | Current Buffer Size | History Buffer Size | Filter Window | Filter |
| X (Shear) | |||||
| Y (Shear) | |||||
| Z (Normal) | |||||
C.3 Training Settings
Simulation Settings. When training the oracle policy, we use 32768 parallel environments, distributed on 4 GPUs, to train the agent. Each environment contains a simulated Allegro Hand, simulated ReSkin palm collision geometries, a cylinder with randomized parameters (Table 6), as well as random initial and goal positions. The ReSkin palm tactile outputs are not simulated for the oracle policy, but we include the collision geometries for the oracle policy, to have consistent dynamics when training the student policy. The simulation frequency is Hz, the control frequency is Hz, each episode lasts for 400 control steps ( s). The wall clock time for oracle policy training is 1.68 days.
When training the tactile and proprioceptive policies, we simulate the ReSkin palm outputs (Table 7), and instead we use 1000 parallel environments distributed across 2 GPUs. Unless mentioned otherwise, the simulation settings are the same as the oracle policy. The wall clock time for training the student policy with tactile skin simulation is 6.27 hours.
Stable Grasp Generation. Both the oracle and tactile policies assume that the object is initialized in a stable grasp, similar to [46, 7]. We pre-generate caches of stable grasp poses for training.
First, we define canonical poses for the fingers, which varies for each scale of the object. Because we train for a large range of scales, the same canonical pose will not work for the entire range. Then, we define a range of relative offsets within rad to randomly sample from. Additionally, we define a canonical position for the object and sample from a relative offset of cm along the x-axis for the object. The canonical object pose is the same for all scales: .
We let the simulation run for 0.5 s. The finger and object positions are saved (i.e., deemed a stable grasp) if the following conditions are satisfied:
1.
All fingertips are within 10 cm of the object.
2.
At least 2 fingers are in contact with the object.
3.
The object did not fall below the hand.
4.
The object is in contact with the palm.
We sampled 50,000 grasps and object poses for each object scale.
Goal Generation. The goal position for the object is sampled from a range of relative offsets cm along the x-axis from the object’s canonical pose. The minimum offset is cm to encourage meaningful object translation, and the maximum is cm which is at the end of the canonical cylinder length.
C.4 Hardware Setup
We place a 0.25mm sheet of PET-like plastic on top of the ReSkin palm to reduce friction between the sensing skin and the object. Reducing the friction will reduce the required torque from the fingers to move the object; this avoids overheating the motors. The plastic sheet is not adhered to the sensing skin.
We also place high-friction 3M tape (3M 300LSE) at the base of thumb and index finger motors. This reduces object slip and somewhat helps with object translation, since the policy sometimes pinches the object between these fingers during part of the finger gait. We also cover the outside casing of the fingertips with the same 3M tape, although they typically do not contact the object.

### v2

@@ VI-B Three-Axis Tactile Sensing Enables Superior Out-of-Domain Adaptation
VI-C Experimental Analysis
Latent Space Analysis. We analyze the extrinsics vector from all 15 rollouts per policy (5 per object) from our experiments in Figure 6 with t-SNE to produce the plots in Figure 7. We find a correlation between high dispersion of clusters and complete policy failure. U3-Axis failed with the hammer at 20 degrees and Proprio Only failed with the water bottle at 5 degrees. The clusters for these failed rollouts are more dispersed compared to more successful rollouts. In contrast, the S3-Axis policy was more successful overall and exhibited a dense cluster of from all rollouts.
Tactile Policies Explore More Finger Gaits. We analyze the joint states of all policies from our experiments in Figure 6. We compare the standard deviation of intersections through Poincaré sections of the phase portraits for each finger motor and find that on average, tactile policies explore more joint states relative to Proprio Only, particularly for the hammer and water bottle. We highlight the contrast in joint exploration in phase portraits of S3-Axis and Proprio-Only, where S3-Axis has an increase of 35% standard deviation of intersections through the Poincaŕe section (Figure 8). Greater joint state exploration is potentially a factor in task success and an indicator of gait adaptation; however, other variables such as gait cycle timing and finger coordination are also important to consider.
VI-D Demonstrations
Diverse Objects. We demonstrate our S3-Axis and U3-Axis policies with an additional 15 objects (Supp. Video). Non-cylindrical objects include the mustard bottle, bleach bottle, and multimeter. The object weights range from g - g. We also demonstrate with varied object compliances: flexible (cheetos, cups), soft (paper towel roll, pool noodle, soft tripod case), and rigid (80-20 aluminum bar, air can).
Real-World Videos. Our Project Website hosts videos of our experiments, which can be used to observe task metrics and policy gait adaptation.
VII Conclusion and Future Work
@@
Our current tactile model is heuristic and empirically formulated to match real ReSkin signals. Future work will explore models more aligned with physics, especially for continuous tactile signals. A key challenge in sim-to-real transfer of ReSkin coverage on both the fingers and palm will be modeling cross-talk between multiple discrete sensors. We can also explore domain adaptation methods for sim-to-real tactile transfer, and investigate how 3-axis tactile sensing improves policy performance for other dexterous skills.
Acknowledgement
@@
We thank Mustafa Mukadam, Mike Lambeta, Tingfan Wu, Luis Pineda, Taosha Fan, Patrick Lancaster, Mrinal Kalakrishnan, Raunaq Bhirangi, Carolina Higuera, Akash Sharma, Suddhu Suresh, and William Yang for helpful discussions and feedback throughout this work. We thank Dr. Nadia Figueroa, Ho Jin Choi, Tianyu Li, and Harshil Parekh for assistance with the Franka and Optitrack system.